Find r so that the line through (2; 6) and (- 4; r) is as follows:

ii. Perpendicular to x + 2y = 1


1
Expert's answer
2021-11-02T16:01:44-0400

"y=1\/2-x\/2"


 perpendicular slope:

"m=2"


equation of perpendicular line:

"y=2x+b"

for  point (2; 6):

"6=4+b"

"b=2"

"y=2x+2"


for  point (- 4; r):

"r=-8+2=-6"